OpenStudy (anonymous):

Charlotte embroidered a piece of cloth with a length of 4 meters and width of 2 meters in 4 1/2 what is the approximate speed in square meters per hour at which Charlotte can embroider?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@Mehek14

OpenStudy (anonymous):

i need steps please!

OpenStudy (boldjon):

pemdas so do parenthaseese first work inside then out (25 times 2) (50) now we have 96/42+(50)-15-3=? start from the begining again of the equation (left) now we do division 96/42=16/7 now we have 16/7+(50)-15-3 start over and do addition 16/7+50=2 and 2/7+50=2+2/7+50=52 and 2/7 subtract 52 and 2/7-15=52+2/7-15=37 +2/7=37 and 2/7 then subtract 3 37 and 2/7-3=37+2/7-3=34+2/7=34 and 2/7 answer is 34 and 2/7

OpenStudy (anonymous):

Thank u!

OpenStudy (boldjon):

where's the medal